OpenStudy (anonymous):

What is the distance from M(-1, 6) to N(11, 1)? A) 12 units C) 13 units B) 149−−−√ units D) 169 units

myininaya (myininaya):

\[Distance=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}\] where \[(x_1,y_1)=(-1,6) ;(x_2,y_2)=(11,1)\]

OpenStudy (anonymous):

immaculate presentation myin

OpenStudy (anonymous):

B)\[\sqrt{149}\]

myininaya (myininaya):

\[\sqrt{(-1-11)^2+(6-1)^2}=\sqrt{(-12)^2+5^2}=\sqrt{144+25}=\sqrt{169}=13\]

OpenStudy (chriss):

It's C. -1 - 11 = -12, -12^2=144 6-1=5, 5^2=25 144+25 is 169 and the square root of 169 is 13.

OpenStudy (anonymous):

i'd do anything for a medal (lol)

OpenStudy (anonymous):

jimmy, are you a teacher?